### v2

# Masked Attribute Description Embedding for Cloth-Changing Person Re-identification

Abstract
Cloth-changing person re-identification (CC-ReID) aims to match persons who change clothes over long periods. The key challenge in CC-ReID is to extract clothing-independent features, such as face, hairstyle, body shape, and gait. Current research mainly focuses on modeling body shape using multi-modal biological features (such as silhouettes and sketches). However, it does not fully leverage the personal description information hidden in the original RGB image. Considering that there are certain attribute descriptions which remain unchanged after the changing of cloth, we propose a Masked Attribute Description Embedding (MADE) method that unifies personal visual appearance and attribute description for CC-ReID. Specifically, handling variable clothing-sensitive information, such as color and type, is challenging for effective modeling. To address this, we mask the clothing and color information in the personal attribute description extracted through an attribute detection model. The masked attribute description is then connected and embedded into Transformer blocks at various levels, fusing it with the low-level to high-level features of the image. This approach compels the model to discard clothing information. Experiments are conducted on several CC-ReID benchmarks, including PRCC, LTCC, Celeb-reID-light, and LaST. Results demonstrate that MADE effectively utilizes attribute description, enhancing cloth-changing person re-identification performance, and compares favorably with state-of-the-art methods. The code is available at https://github.com/moon-wh/MADE.
1 Introduction
Person re-identification (Re-ID) refers to the process of identifying persons across different camera views. Traditional Re-ID methods operate under the assumption that the person being tracked moves within a confined area and time and will not change their clothes. However, in practical scenarios, persons captured by surveillance cameras may traverse larger areas and be observed over extended periods, during which they might change their clothing. This deviation in appearance challenges the reliability of color-based information utilized by earlier Re-ID approaches for person re-identification. Therefore, in recent years, cloth-changing person re-identification has attracted more and more attention.
In cloth-changing person re-identification, the reliance on color information by traditional methods becomes unreliable, and addressing the modeling of clothing changes poses a significant challenge [44]. Therefore, the key to solving CC-ReID lies in identifying information about personal clothing that is insensitive to these changes. To mitigate the interference caused by varying clothes and uncover invariant features of persons, some cloth-changing person re-identification methods focus on the multi-modal biometric features of persons. PRCC [44] and FSAM [19] study silhouette information of persons. 3DSL [4] aims to learn the 3D shape features of persons. However, contours and 3D features eliminate all color information of images. MBUNet [47] contains a branch for extracting posture features. GI-ReID [23] and ViT-VIBE [2] utilize gait features. However, extracting pose features from a single image of a person is still a challenging task. SpTSkM [48] uses skeleton normalization to assist person recognition. RF-ReID [11] infers personal skeleton features from radio frequency signals. However, skeleton features are challenging to extract and utilize directly. These methods often require additional complex models for extracting multi-modal biometric features, demanding substantial computing resources for training. This complexity overlooks the most direct description information present in personal images. Moreover, the way these methods utilize the extracted clothing-insensitive information is usually more complex. CAL [15] and the method proposed by chan et al. [3] are both based on GAN [14] network.
Figure 1: In the scene of cloth-changing person re-identification, the person captured across cameras has many attributes that remain unchanged (attribute words marked in blue) except for their clothes (attribute words marked in black).
In the context of cloth-changing person re-identification, even over extended time intervals, persons tend to alter only their clothing choices, while other attributes such as gender, age, and hair color remain consistent, as depicted in Fig. 1. These clothes-interrelated attributes are helpful to person re-identification, while clothes-related attributes can be easily eliminated. This paper introduces the Masked Attribute Description Embedding (MADE) method to effectively mine clothing-insensitive information from original RGB images of persons. Specifically, we adopt the attribute detection model to extract pedestrian attributes. Then, the clothing and color information is masked to obtain masked attribute description, and the clothing-sensitive features are eliminated by shielding the clothing information in the RGB image and retaining other clothing-insensitive color features. Due to the editable nature of text description, clothing color can be quickly and efficiently eliminated. MADE then connects and embeds the masked attributes description data encoded by Linear Projection into TrV blocks [12] at different levels to fuse the image features. By mapping different feature spaces to a shared latent space, masked description can be fused with image features without the need for additional text encoders, forcing the model to discard clothing-sensitive information.
We summarise the contributions of this work as follows.
  • We propose a Masked Attribute Description Embedding (MADE) re-identification method for cloth-changing person re-identification, which unifies the person’s variable color visual appearance and editable attribute description in CC-ReID.
  • We introduce multi-modal attribute description information in CC-ReID that is more evident and easier to extract and edit than skeletons or contours in the original person’s images. Since the clothing-sensitive information is not conducive to identifying persons, we use the attribute detection model to extract the description in Description Extraction and Mask(DEM) module, mask the variable clothing and color items and embed it into the image features, which can force the model to discard clothing-sensitive features.
  • For the first time, we employ a simple and efficient method to integrate image features with attribute description in CC-ReID. MADE connects and embeds the masked attribute description data, encoded by Linear Projection, into transformer blocks at various levels, facilitating the fusion of low-level and high-level image features. Through the mapping of different feature spaces to a shared latent space, the model gains a deeper understanding and exploitation of semantic relationships. Without the need for additional text encoders, descriptions seamlessly fuse with image features, enabling the model to capture the associated information between images and descriptions effectively.
  • Our extensive experiments on four public benchmark datasets, PRCC, LTCC, Celeb-ReID-light, and LaST, show that MEDA consistently outperforms existing state-of-the-art methods by a large margin.
In the following, we will discuss related work in section 2. We will present the details of our poposed method in section 3. The experimental results are provided in section 4. Section 5 concludes this paper with our future reseach directions
2 Related Work
2.1 Cloth-Changing Person Re-Identification
The core of solving cloth-changing person re-identification is to extract the clothes-irrelevant features in person images. To this end, some research focuses on multi-modal biometric features that are less variable than clothing, such as silhouette, 3D shape, skeleton, walking posture, etc.
FSAM [19] uses a parsing network to train and obtain contour images, enabling coarse-to-fine mask learning. [5] proposes a multi-scale appearance and contour depth Infomax (MAC-DIM) to maximize the mutual information between appearance and contour shape features. Mu et al. [30] utilize human parsing models to segment the semantic parts of the human body to obtain the binary body shape masks. These methods discard all color information in the original RGB image in the contour processing module. However, some color information is helpful for person re-identification.
SpTSkM [48] explores persons’ motion pattern information from 3D skeletons normalized by ST-GAN [43] to assist person re-identification. 3DSL [4] distinguishes different identities by learning 3D shape features and 3D reconstruction subnetworks. These methods obtain 3D shapes through cumbersome 3D parsing and processing networks, which increases the complexity of model training.
CESD [31] uses a pose detector to detect personal body joint points and uses shape embedding to separate clothing and distinguish shape information through joint point features. The pose feature branch of MBUNet  [47] applies the direction adaptive graph convolution layer to obtain the relevant information between different keypoints in heatmaps. ViT-VIBE [2] uses ViT [10] to combine appearance and gait features learned through VIBE [25]. However, these methods do not fully exploit and utilize the clothing-independent features in the original image.
Some methods use adversarial strategies to mine clothing-insensitive information in original images. AFD-Net [42] produces more realistic images by using a discriminator for data augmentation. CCFA [16] uses an adversarial loss to supervise the augmentation vector to ensure that it roughly satisfies the empirical clothing change distribution in the training data. CAL [15] classifies clothing based on a GAN network and penalizes the model’s ability to predict clothing. However, adversarial models are prone to training instability and have difficulty explaining how to separate appearance and clothing information from RGB images.
In this paper, we propose MADE to unify persons’ appearance and language description. Multi-modal attribute description information is introduced in CC-ReID, which is more obvious and accessible to extract and edit than biological features such as skeleton or silhouette in original person images. Information that helps identify persons can be retained to the greatest extent while accurately removing interference from clothing information.
2.2 Text-to-Image Person Retrieval
Li et al[27] created the first benchmark dataset CUHK-PEDES. The main challenge in this field is how to efficiently fuse image and text features into a joint embedding space. Early research work [26, 27] adopted VGG [38] and LSTM [18] to learn the representation of visual-text modalities. Subsequent studies [7, 35, 49] designed new cross-modal matching losses to align global image-text features in the joint embedding space. IRRA [22] integrates visual cues with CLIP [33] encoded text tokens into a cross-modal multi-modal interaction encoder, enabling cross-modal interaction. These methods utilize sentence captions describing persons. However, when dealing with scenes involving changes in clothing, the model faces challenges in directly processing clothing-related fragments within the captions. Introducing an encoder to encode language text would increase the complexity of model training. To address this, we leverage itemized attributes. This approach enables the model to precisely handle clothing-related information within the description, avoiding the processing of the entire statement as a whole. Simultaneously, we embed the attribute vector directly into the Transformer block, eliminating the need for additional encoder encoding.
2.3 Attribute based Person Re-identification
It has been well exploited to perfrom person re-identification with attributes. APR [28] introduces the Attribute Reweighting Module (ARM), which corrects predictions of attributes based on learned dependencies and correlations between attributes. MSPA [24] uses ConvLSTM to memorize the relationship between persons’ attribute features. These methods utilize all pedestrian attributes to address traditional person re-identification challenges but overlook the editability of the description. We introduce the editability of description in CC-ReID, which can accurately remove clothing-sensitive attributes and help the model learn clothing-sensitive information.
3 Method
3.1 Overview
The key to achieving cloth-changing person re-identification lies in extracting clothing-insensitive features from the image. In CC-ReID, we introduce attribute description to mitigate the impact of clothing interference. Consequently, our model primarily focuses on modeling the relationship between and within the two types of images and descriptions. EVA-02 [12] is pre-trained to reconstruct powerful and robust language-aligned visual features through occlusion image modeling, resulting in transferable models. Based on this model, we proposed MADE framework to integrate persons’ masked attribute description data with image visual features, addressing the challenges of cloth-changing person re-identification. The framework of our approach is shown in Fig. 2. Given an image sample $x_{i}$, $x_{i}$ extracts an editable attribute description through Description Extraction and Mask module. After the masked attribute description is converted into a binary vector, it is connected and embedded at different levels through Linear Projection to fuse with image features in TrV blocks [12]. We introduce how to extract and mask attribute description in Sec. 3.2. Then, Sec. 3.3 details how to add masked attribute description data to improve the performance of CC-ReID. Finally, we elaborate on the model’s loss function and inference process.
Figure 2: The framework of Masked Attribute Description Embedding (MADE) method. We first extract editable attribute description from the image through Description Extraction and Mask (DEM) module. After the cloth-related attribute descriptions are masked and converted into a binary vector, it is connected and embedded at different levels through Linear Projection to fuse with image features. Finally, we aggregate $f_{cls}^{v}$, $f_{des,2}^{m}$ and $f_{des,3}^{m}$ through Conv1D to obtain the person feature representation
3.2 Description Extraction and Mask (DEM)
In CC-ReID, it is imperative for the model to disregard clothing information in the RGB image during input to learn clothing-insensitive features. DeSKPro [41] and SAVS [13] use the human parsing model to generate person parsing maps to remove clothing interference. However, processing the parsing map to obtain robust cloth-irrelevant information is more complicated. The description information of persons mainly describes the appearance and clothing of persons, so previous research rarely involves processing person description in CC-ReID. However, we could eliminate interference information from person’s clothes in the model input by simply editing the attribute description. In MADE, for each input image $x_{i}$, we extract the description and operate the clothes mask through DEM, as shown in Fig. 2.
DEM uses the extraction model to obtain persons’ descriptions suitable for cloth-changing datasets and performing mask processing. Specifically, we use SOLIDER [6] trained on PETA [9] to identify person attributes in cloth-changing datasets. SOLIDER [6] is a human task visual pre-training model that adopts self-supervised training. We use it to obtain attribute description of images. PETA contains 19,000 images, including 8,705 individuals, each annotated with 61 binary and four multiclass attributes, 105 attribute labels in total. Given a sample $x_{i}$, input SOLIDER to get persons’ attributes list, [$a_{1}^{i}$, $a_{2}^{i}$,…,$a_{N}^{i}$], and convert them into 0-1 binary vectors by position. We implement the clothing information mask operation by setting the clothing attributes to 0. And, the masked attribute description data [$m_{1}^{i}$, $m_{2}^{i}$,…,$m_{N}^{i}$] with the clothing information removed is obtained.
In all experiments, we train the model with parameter settings that follow the original project [6].
3.3 Process of Description Embedding
More than relying on appearance information is required to distinguish persons who change clothes accurately. Immutable multi-modal features can assist recognition [19, 31]. The advancement of Text-to-Image Person Retrieval demonstrates that additional text information can be fully leveraged when learning images of persons to enhance the final decision-making process. However, existing approaches often utilize a text encoder to handle entire sentences. We propose MADE to embed the masked attribute description vector directly into the Transformer block.
Given an image sample $x_{i}\in R^{H\times W\times C}$and its masked-description data $m_{i}$ = [$m_{1}^{i}$, $m_{2}^{i}$,…,$m_{N}^{i}$], we integrate them into TrV Block [12] to get the MADE framework, as shown in Figure 2. TrV Block is an improved Vision Transformer structure. We use EVA02-large as the backbone, which has 24 layers of TrV Blocks and divides it into three stages (we discuss it in detail in 4.6.2). First, we segment the image $x_{i}$ into a sequence of $N=H\times W/P^{2}$ fixed-size, where P represents the size of the patch, and then map the patch sequence through a trainable linear projection as one-dimensional notation {${f_{j}^{v}|_{j=1}^{N}}$}. After the injection of positional embedding and additional [CLS] token, the tokens sequence {$f_{cls}^{v},f_{1}^{v},…,f_{N}^{v}$} is input into the TrV block of $L$ layers of the first stage to model dependencies between each patch. Subsequently, $f_{cls}^{v}$ is extracted to represent the first stage’s global image low-level feature representation. Then $m_{i}$ is passed through Linear Projection to obtain the description feature {${f_{j}^{m}|_{j=1}^{N}}$} with the same dimension as {${f_{j}^{v}|_{j=1}^{N}}$}. After adding the extra token [DES], the attribute description sequence is {$f_{des}^{m},f_{1}^{m},…,f_{N}^{m}$}. The description sequence is then connected to the image sequence, {$f_{j}^{vm}|_{j=1}^{N}$}={$f_{des}^{m},f_{1}^{m},…,f_{N}^{m},f_{1}^{v},…,f_{N}^{v}$}, input the second and third stages for training. In this process, image features and description features are embedded through connections and trained together to learn the relationship between images and attribute description that mask clothing information. The interference of clothing-sensitive can be removed through mask items and connection embedding, avoiding the problem of complex extraction and fusion of multi-modal biometric features. In the model, we extract $f_{des}^{m}$ as a fusion representation of image and attribute description features. The class tokens output by the second and third stages, $f_{des,2}^{m}$ and $f_{des,3}^{m}$, respectively, represent the fusion of different levels of visual features and attribute description of clothing removal. We combine them with the output of $f_{cls}^{v}$ from the first stage to obtain the pedestrian feature representation of MADE through Conv1d aggregation.
3.4 Loss Function and Inference
In our experiments, we use cross-entropy loss without label smoothing and triplet loss. the loss function of MADE can be defined as follows:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{id}+\lambda_{2}\mathcal{L}_{tri}$ (1)
Where $\mathcal{L}$ is the total loss function of the MADE method, $\mathcal{L}_{id}$ represents cross-entropy loss, and $\mathcal{L}_{tri}$ represents triplet loss. $\lambda_{1}$ and $\lambda_{2}$ are trade-off parameters used to balance each contribution. In our experiments, both $\lambda_{1}$ and $\lambda_{2}$ are empirically set to 1.0.
Cross-entropy loss $\mathcal{L}_{id}$ is defined as:
$\mathcal{L}_{id}=-\sum_{i=1}^{N}\log\frac{\exp(W_{y_{i}}x_{p}^{i}+b_{y_{i}})}{\sum_{k=1}^{C}\exp(W_{k}x_{p}^{i}+b_{k})}$ (2)
where $N$ is the number of images in mini-batch, $y_{i}$ is the label of feature $x_{p}^{i}$ and $C$ is the number of classes.
Triplet loss function $\mathcal{L}_{tri}$ is defined as follows:
$\mathcal{L}_{tri}(I_{iA},I_{iP},I_{iN})=max\{0,M+D(I_{iA},I_{iP})-D(I_{iA},I_{iN})\}$ (3)
where $D(.)$ is the squared Euclidean distance in the embedding space, and $M$ is a parameter called the margin that adjusts the separation between pairs of distances: ($f_{iA},f_{iP}$) and ($f_{iA},f_{iN}$). $I_{iA}$, $I_{iP}$, and $I_{iN}$ are anchor images, positive samples, and negative sample images, respectively. The model learns to minimize the distance between more similar images and maximize the distance between dissimilar images.
For inference, for a given query $q_{i}$ and masked attribute description data $m_{i}$ = [$m_{1}^{i}$, $m_{2}^{i}$,…,$m_{N}^{i}$], we only use query images $q_{i}$ and discard $m_{i}$ to inference.
4 Experiments
4.1 Datasets
PRCC [44] is a dataset including person contour sketch proposed by Yang et al., including 221 persons and 33,698 images. The photos are taken by three cameras, A, B, and C, respectively, and the clothes of persons in cameras A and B do not change. Camera C takes pictures at different times, and the clothes of persons are different from those in cameras A and B. There are about 50 images of each person under each camera view.
LTCC [31] is a dataset captured by 12 cameras for two months, including 162 persons and 15,138 images. The dataset is divided into two subsets: persons with changing clothes, including 91 people, 415 sets of different clothes, and 14,756 images; the set with consistent clothes, including 61 people and 2,382 images.
Celeb-reID-light [20] contains 290 persons and 10,842 images. This dataset comes from the snapshots of celebrities on the Internet. Everyone in the dataset has about 20 pictures of different clothes, and people do not wear the same clothes.
LaST [37] is a large-scale dataset from over 2,000 movies in 8 countries. It includes 10,862 persons and 228,166 images. The training set has 5,000 identities and 71,248 images, the validation set has 56 identities and 21,379 images, and the test set has 5,806 identities and 135,529 images.
4.2 Attribute Analysis
In order to explore whether the irrelevant attributes of clothes are retained when persons change clothes and are captured across cameras, and the proportion of retention. In the experiment, we evaluated the retention ratios of attributes in the training and test sets for the four datasets: PRCC, LTCC, Celeb-ReID-light, and LaST.
As depicted in Fig. 3, the figure illustrates the higher maintained proportions of cloth-irrelevant attributes in each dataset. It is observed that the biological attributes (age, gender, hair) and carrying objects of persons generally do not change across cameras in the short term. Additionally, features like shoe type and color generally stabilize even as persons change their clothing. These attributes can be important in the model learning clothing-insensitive features.
Refer to caption (a) PRCC Refer to caption (b) LTCC Refer to caption (c) Celeb-reID-light Refer to caption (d) LaST
Figure 3: Retain ratio of irrelevant attributes of persons’ clothing in each dataset.
4.3 Evaluation Settings and Protocol
For the cloth-changing datasets PRCC [44] and LTCC [31], we follow their respective evaluation protocols and evaluate the performance under the clothes-changing and standard settings.
We adopt the standard metrics used in most of the person re-identification literature, namely the cumulative matching curve (CMC), to generate ranking accuracy and the mean average precision (mAP). We report rank-1 accuracy and mean average precision (mAP) on all datasets for evaluation.
4.4 Implementation Details
The input images are resized to 224×224 for all datasets. We divide the 24 layers of EVA02-large  [12] into three stages, and the number of layers of the Trv blocks is 10, 10, and 4, respectively(We discussed in 4.6.2). We use LayerNorm  [1] to normalize features. For data augmentation, we employ random cropping and random erasing [51]. Due to the limit of GPU memory, the batchsize is set to 8, each batch includes two different people, and the number of images for each person is 4. The SGD optimizer is employed in the optimization process, and 60 epochs are required. Moreover, the weight decay for the experiment is $5e^{-2}$. The warmup learning rate is initially set to $7.8125e^{-7}$. The learning rate is initially set to $2e^{-5}$ and divided by 100 at 40 and 60 epochs. The optimal parameter values are directly used for the other datasets without tuning.
4.5 Comparison with State-of-the-Art
For RRCC and LTCC, we combine our proposed MADE method with some cloth-changing re-identification methods (i.e., SPT+ASE [44], GI-ReID [23], CESD [31], RCSANet [21], 3DSL [4], FSAM [19], BSGA+CRE [30], CAL [15], DCR-ReID [8], CCFA [16], AIM [45] and chan et al. [3]) were compared. We compare Celeb-reID-light with four cloth-changing re-identification methods (RCSANet [21], MBUNet [47], IRANet [36], and DeSKPro [41]) and some traditional methods. We compare LaST with CAL and some traditional methods. It is worth noting that among these CC-ReID methods, SPT+ASE, GI-ReID, CESD, 3DSL, FSAM, BSGA+CRE, and DCR-ReID all integrate person multi-modal biometric features into the model to remove clothing interference. CAL and AIM mine the information of original RGB images. CCFA adopts the feature enhancement method. The method proposed by Chen et al. is based on the GAN network.
Table 1: Evaluations on the PRCC and LTCC datasets (%), where ”sketch,” ”pose,” ”sil.”, ”parsing” and ”3D” denote contour sketches, keypoints, silhouettes, human parsing, and 3D shape information. Bold and underlined numbers are the top two scores.
Method Venue Modality PRCC LTCC CC SC CC Genral rank-1 mAP rank-1 mAP rank-1 mAP rank-1 mAP SPT+ASE  [44] TPAMI 19 Sketch 34.4 - 64.2 - - - - - CESD  [31] ACCV 20 RGB+pose - - - - 26.1 12.4 71.4 34.3 RCSANet  [21] ICCV 21 RGB 48.6 50.2 100.0 97.2 - - - - 3DSL  [4] CVPR 21 RGB+pose+sil.+3D - 51.3 - - 31.2 14.8 - - FSAM  [19] CVPR 21 RGB+pose+sil. 54.5 - 98.8 - 38.5 16.2 73.2 35.4 GI-ReID  [23] CVPR 22 RGB+sil. - 37.5 - - 23.7 10.4 63.2 29.4 BSGA+CRE  [30] BMVC 22 RGB+parsing 61.8 58.7 99.6 97.3 - - - - CAL  [15] CVPR 22 RGB 55.2 55.8 100.0 99.8 40.1 18.0 74.2 40.8 CCFA  [16] CVPR 23 RGB 61.2 58.4 99.6 98.7 45.3 22.1 75.8 42.5 AIM  [45] CVPR 23 RGB 57.9 58.3 100.0 99.9 40.6 19.1 76.3 41.1 DCR-ReID  [8] TCSVT 23 RGB+parsing 57.2 57.4 100.0 99.7 41.1 20.4 76.1 42.3 chan et al. [3] ACM 23 RGB 58.4 58.6 100.0 99.7 32.9 15.4 73.4 36.9 MADE RGB+description 67.5 64.2 100.0 98.9 46.9 25.0 82.2 49.3
Results on PRCC. We compare our method with twelve cloth-changing re-identification methods on PRCC in Table 1. We can notice that our method outperforms all other methods in the clothes-changing setting. Compared with AIM, the method of mining original RGB images in the clothes-changing setting, the rank-1 of our method increased by 9.6%, and the mAP increased by 5.9%. These shows that using multi-modal attribute description in CC-ReID to assist re-identification effectively improves results. Compared with the best method using multi-modal biometrics, BSGA+CRE, in the clothes-changing setting, the rank-1 increased by 5.7%, and the mAP increased by 5.5%. It shows that our method uses editable multi-modal attribute information and has better results when it is more convenient to remove clothing information than biological information. Data augmentation can significantly improve the model improvement effect, and our method is even better than CCFA, which uses feature enhancement. In the clothes-changing setting, the rank-1 increased by 6.3%, and the mAP increased by 5.8%.
Results on LTCC. We compare our method with eight cloth-changing re-identification methods on LTCC in Table 1. Our method outperforms all other methods in both clothes-changing and general settings. Compared with CCFA, in the clothes-changing setting, The rank-1 increased by 1.6%, and the mAP increased by 2.9%. In the general setting, The rank-1 increased by 6.4%, and the mAP increased by 6.8%. Compared with AIM, in the clothes-changing setting, The rank-1 increased by 6.3%, and the mAP increased by 5.9%. In the general setting, The rank-1 increased by 5.9%, and the mAP increased by 8.2%. Compared with chan et al., in the clothes-changing setting, The rank-1 increased by 14.0%, and the mAP increased by 9.6%. In the general setting, The rank-1 increased by 8.8%, and the mAP increased by 12.4%. It shows that our method embedding fused image features with attribute description is better than using GAN networks to mine clothing-independent features of original images.
Table 2: Evaluations on Celeb-reID-light(%). Bold and underlined numbers are the top two scores.
Method Type Method Venue Celeb-reID-light rank-1 mAP Traditional OSNet [52] ICCV 19 21.3 11.7 DG-Net [50] CVPR 19 23.5 12.6 BoT(resnet50) [29] CVPRW 19 24.2 13.6 AGW(resnet50_nl) [46] TPAMI 21 30.2 15.4 TransReID [17] ICCV 21 31.3 18.6 CC-ReID RCSANet  [21] CVPR 21 29.5 16.7 MBUNet  [47] ICME 22 33.9 21.3 IRANet  [36] IVC 22 46.2 25.4 DeSKPro  [41] ICIP 22 52.0 29.8 MADE 69.6 52.1
Table 3: Evaluations on LaST(%). Bold and underlined numbers are the top two scores.
Method Type Method Venue Celeb-reID-light rank-1 mAP Traditional OSNet [52] ICCV 19 64.3 21.0 BoT [29] CVPRW 19 67.1 23.6 HOReID [40] CVPR 20 68.3 25.5 Top-DB-Net [32] ICPR 20 69.4 25.0 CtF [39] ECCV 20 70.0 26.5 CC-ReID CAL [15] CVPR 22 73.7 28.8 MADE 82.6 45.8
Results on Celeb-reID-light and LaST. We compare our method with some traditional and cloth-changing re-identification methods on the two datasets in Table 2 and Table 3. Our method outperforms all previous methods. The face is the most direct information for re-identification. For Celeb-reID-light, our method is even better than DeSKPro, which uses facial features. The rank-1 increased by 17.6%, and the mAP increased by 22.3%. For LaST, our method is better than CAL. The rank-1 increased by 8.9%, and the mAP increased by 17.0%. LaST is a large and challenging dataset that requires high model training complexity. Currently, there are few CC-ReID methods tested using LaST. It shows that our method is effective, has low model complexity, and can achieve testing of large datasets.
4.6 Ablation Study
We use EVA02-large  [12] encoded by cloth label embedding  [34] as the baseline and also use the loss function in 3.4 for supervision. In this section, we explore the role of masked attribute description on the model’s learning of clothes-irrelevant features and the impact of different layer numbers in EVA02-large.
4.6.1 Effectiveness of Description
We embed the complete attribute description vector (attribute description) and the attribute description vector after mask clothing information (masked attribute description) into the baseline, respectively, and summarize the experimental results in table4. To verify the effectiveness of attribute description in improving re-identification results. We conducted experiments on the clothes-changing setting of PRCC and LTCC. In both datasets, the performance of adding attribute description and masked attribute description is almost higher than the baseline.
The results of adding attribute description are higher than the baseline, indicating that even after persons change clothes, clothing-related attributes still play an essential role in the model. After masking the clothing information in the attribute description and embedding, for PRCC, the Rank-1 increased by 5.0%, and the mAP increased by 5.9%. For LTCC, the Rank-1 increased by 0.7%, and the mAP increased by 1.1%. The model is more accurate in person re-identification, indicating that this method can compel the model to discard clothing information and learn clothing-insensitive features.
Table 4: Ablation studies of attribute description of MADE in clothes-changing setting on PRCC and LTCC.
Method PRCC LTCC rank-1 mAP rank-1 mAP baseline 62.5 58.3 46.2 23.9 baseline + attribute 66.3 62.7 45.4 24.2 baseline + masked attribute(MADE) 67.5 64.2 46.9 25.0
4.6.2 Number of Stages and TrV Blocks
In this section, we discuss the layering situation of EVA02-large, conduct experiments with MADE, and fuse attribute description features that remove clothing information with low-level to high-level features of images. We tried several stratification scenarios based on experience and summarized the experimental results on the clothes-changing setting of PRCC in Table 5. We can observe that the best results are achieved when the number of stages is three, and the number of layers is 10, 10, and 4, respectively.
Table 5: Ablation studies of the different number of stages and TrV blocks for MADE in clothes-changing setting on PRCC. Bold numbers the top score.
# Stage # Layer (24) PRCC rank-1 mAP 2 [12, 12] 57.6 51.5 2 [10, 14] 48.3 45.7 3 [8, 8, 8] 63.7 57.6 3 [8, 12, 4] 64.4 60.0 3 [10, 10, 4] 67.1 62.7 4 [6, 6, 6, 6] 54.0 49.0 4 [4, 4, 8, 8] 54.5 51.7
5 Conclusion
We propose the Masked Attribute Description Embedding (MADE) method, which integrates a person’s visual appearance with attribute description in CC-ReID. The modeling of volatile clothing-sensitive information, including color and type, is challenging and not conducive to identifying persons in CC-ReID. To address this, we introduce multi-modal attribute description information in CC-ReID, which is more obvious and easier to extract and edit than skeletons or contours in original images. We extract descriptions suitable for persons’ images using an attribute detection model, mask the variable clothing and color information, and embed it into the image features, compelling the model to discard clothing information. Subsequently, MADE connects and embeds the masked attribute description features encoded by Linear Projection into Transformer blocks at different levels, fusing them with low-level to high-level features of the image. By mapping different feature spaces to a shared latent space, attribute description can be fused with image features, enabling the model to capture the associated information between images and descriptions effectively. We conducted experiments on PRCC, LTCC, Celeb-reID-light, and LaST. Extensive experiments have demonstrated that MADE can effectively utilize persons’ description information to improve the performance of cloth-changing person re-identification and performs well compared to state-of-the-art methods.
In the future, we intend to expliot the Large Language Models (LLM) to generate better attribute descriptions, which could help futher improve the generalization ability of our method for cloth-changing person re-identification. We will also explore the possibility of our masked attribute description strategy in other cross-modality person re-identification tasks, such as visible-infrared person Re-ID, cross-resolution person Re-ID, and sketch based person Re-ID, because there are also certain attributes which remain unchanged across the modalities.